2.5 2#7 fc= 4,000 psi fy= 60,000 psi n 8 5#11 3 18 For service level moment M= 200 ft-k, use the procedures from Chapter 2
Answer #1

la) Lorating neutras asin 028 Mz200 ft-k 18* X + (24-1) 1-2 (2-2.5 ) = 8x7.8 x (21-X) 90 + 27:62-69 = 1310.4 - 62.42 (2n-1) 1(C) Astumins fy = ty Asdy a 0.85 fb B, C + nády 27 (12 in ) 18:57 (7 ) r 78 x 60 = 0-85*4 x 18x085 C + 12x66 C = 7.61 is a =